# A Fluid-Structure Interaction Model of the Zebrafish Aortic Valve

### v1

@@ 4 Discussion and Conclusions
The leaflet kinematics also showed similarities, with some notable contrasts, to human studies. The qualitative kinematics were strikingly similar to those we observed in previous simulations, with similar appearance of the open configuration and orifice, as well as the closed configuration including appearance of slight excess free edge tissue (Kaiser et al.; 2024). One notable difference was the closing transient, which showed a single negative oscillation, rather than a sustained oscillation with multiple sign changes in flow rate, that appeared proportionally longer in duration than that we observed in previous human simulation studies. It has been hypothesized and shown that vortices shed from the leaflet and in the aortic sinus assist in human aortic valve closure (Bellhouse and Talbot; 1969). Given the Reynolds number and lack of vortices or a sinus, this mechanism does not occur in our simulations of the zebrafish valve, but the valve closes without leak nonetheless. This work thus suggests that while vortices may assist closure, this mechanism is not necessary for valve closure. The vortex, however, may shorten the closing transient and reduce associated regurgitant flow.
Examining leaflet mechanics, experimental estimates on human valves predicted a tangent modulus of $9.9\pm 1.8\cdot 10^{7}$ dynes/cm2 circumferentially and $2.3\pm 0.4\cdot 10^{7}$ dynes/cm2 radially (Pham et al.; 2017). Thus, the predicted tangent moduli in the zebrafish are two orders of magnitude lower that that of humans. Other studies were qualitatively similar to our results. A combined experimental/simulation study showed elevated stress at the commissures and somewhat elevated stress in the leaflet belly (Balguid et al.; 2008). Another combined experimental/simulation study showed elevated radial stretch near the annulus, decreasing into the leaflet belly and lower still near the free edge, and elevated circumferential stretch and in the leaflet belly (Rego et al.; 2022).
@@
This study has limitations, in particular we lack experimental measurements of material properties and in vivo valve geometry to validate our predictions. The test chamber is a cylindrical shape, rather than a beating heart or anatomical shape due to lack of in vivo images at this developmental timepoint. We relied on an ex vivo image of the zebrafish heart that was collapsed without pressure for estimating annular radius and height.
@@ 5 Acknowledgements
All authors were supported in part by NIH grants 5R01HL129727-08 and 5R01HL159970-04. EZ was supported in part by NIH grant T32HL144449 Computing for this project was performed on the Stanford University’s Sherlock cluster with assistance from the Stanford Research Computing Center. We would like to thank David Traver (UCSD) and Nathan Lawson (University of Massachusetts Medical School) for providing the fish line.

### v2

@@ 4 Discussion and Conclusions
This work demonstrates the effectiveness and validity of the design-based elasticity concept. In the zebrafish, velocities, pressure and length scales are very different from those used in previous humans studies. Despite these differences, the valve’s physiological function suggests that the estimated stresses and tangent moduli are realistic. That design-based elasticity could be applied to the zebrafish, in the absence of experimental data that would typically be required to conduct mechanics simulations, demonstrates that it is a powerful tool for modeling in biomechanics. These methods are especially useful in the presence of limited data and potentially applicable to a broad set of problems.
@@
The leaflet kinematics also showed similarities, with some notable contrasts, to human studies. The qualitative kinematics were strikingly similar to those we observed in previous simulations, with similar appearance of the open configuration and orifice, as well as the closed configuration including appearance of slight excess free edge tissue (Kaiser et al.; 2024).
One notable difference was the closing transient, which showed a single negative oscillation, rather than a sustained oscillation with multiple sign changes in flow rate, that appeared proportionally longer in duration than that we observed in previous human simulation studies. Doppler echocardiography has detected outflow tract valve regurgitation in some adult zebrafish, both in wild-type individuals and in alk5a-/- mutants (Bensimon-Brito et al.; 2022). Qualitatively similar closing transients in flow rate without regurgitation to those observed in this work have been observed in studies on dogs (Laniado et al.; 1976). It has been hypothesized and shown that vortices shed from the leaflet and in the aortic sinus assist in human aortic valve closure (Bellhouse and Talbot; 1969). Given the Reynolds number and lack of vortices or a sinus, this mechanism does not occur in our simulations of the zebrafish aortic valve, but the valve closes without leak nonetheless. This work thus suggests that while vortices may assist closure, this mechanism is not necessary for valve closure. The vortex, however, may shorten the closing transient or lower its magnitude.
Examining leaflet mechanics, experimental estimates on human valves predicted a tangent modulus of $9.9\pm 1.8\cdot 10^{7}$ dynes/cm2 circumferentially and $2.3\pm 0.4\cdot 10^{7}$ dynes/cm2 radially (Pham et al.; 2017). Thus, the predicted tangent moduli in the zebrafish are two orders of magnitude lower that of humans. Other studies showed qualitatively similar trends to our results. A combined experimental/simulation study showed elevated stress at the commissures and somewhat elevated stress in the leaflet belly (Balguid et al.; 2008). Another combined experimental/simulation study showed elevated radial stretch near the annulus, decreasing into the leaflet belly and lower still near the free edge, and elevated circumferential stretch and in the leaflet belly (Rego et al.; 2022).
@@
This study has limitations, in particular we lack experimental measurements of material properties and in vivo valve geometry to validate our predictions. The test chamber is a cylindrical shape, rather than a beating heart or anatomical shape due to lack of in vivo images at this developmental timepoint. We relied on an ex vivo image of the zebrafish heart that was collapsed without pressure for estimating annular radius and height. In future work, in vivo geometry of the valve and outflow tract could be incorporated as could heterogeneous or three-dimensional, non-membranous thickness of the valve leaflets. The application of similar methods to thick structures that cannot be viewed as membranous or structures with heterogeneous thickness is considerably more complex, due to equations (1) becoming three dimensional and likely having a more complex pattern of strain.
@@ 5 Acknowledgements
All authors were supported in part by NIH grants 5R01HL129727-08 and 5R01HL159970-04. EZ was supported in part by NIH grant T32HL144449. Computing for this project was performed on the Stanford University’s Sherlock cluster with assistance from the Stanford Research Computing Center. We would like to thank David Traver (UCSD) and Nathan Lawson (University of Massachusetts Medical School) for providing the fish line.